Solve the simultaneous equations: y=2x,2y-2x=2
Answer by Cromlix(4381)

y = 2x
2y - 2x = 2
Substitute 2x for y
2(2x) - 2x = 2
4x - 2x = 2
2x = 2
x = 1
Substitute x = 1 in
y = 2x
y = 2(1)
y = 2
x = 1 and y = 2
Hope this helps.
:-)